Quantitative Problem 1: You plan to deposit $2,500 per year for 5 years into a money market account with an annual return of 2%. You plan to make your first deposit one year from today. Do not round intermediate calculations. Round your answers to the nearest cent.

What amount will be in your account at the end of 5 years? $

Assume that your deposits will begin today.

What amount will be in your account after 5 years? $ Quantitative Problem 2: You and your wife are making plans for retirement. You plan on living 30 years after you retire and would like to have $100,000 annually on which to live. Your first withdrawal will be made one year after you retire and you anticipate that your retirement account will earn 10% annually. Do not round intermediate calculations. Round your answers to the nearest cent. What amount do you need in your retirement account the day you retire? $

Assume that your first withdrawal will be made the day you retire. Under this assumption, what amount do you now need in your retirement account the day you retire
Answer #1

1.
=2500/2%*(1.02^5-1)=13010.1004

2.
=2500/2%*(1.02^5-1)*1.02=13270.302408

3.
=100000/10%*(1-1/1.1^30)=942691.446698832

4.
=100000/10%*(1-1/1.1^30)*1.1=1036960.59136872